OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4? y = 4x - 1 y = 4x - 5 y = 4x - 10 y = 4x + 1

jimthompson5910 (jim_thompson5910):

every answer choice has a slope of 4, so all you need to do is plug in x = 3 (from the given point) and if you get y = 2, then that's your answer for instance, y = 4x - 1 y = 4*3 - 1 ... plug in x = 3 y = 12 - 1 y = 11 since this is NOT y = 2 like we want, we can eliminate this answer choice

OpenStudy (anonymous):

man its too easy.. if the point(3,2) is on the line it must satisfy the equation of line .. as in options all lines have slope 4 check each line by putting x=3 y =2

OpenStudy (anonymous):

im still confused. so whats the answer? because what the multiple choice answers are dont say "y=2"

jimthompson5910 (jim_thompson5910):

what happens when you plug x = 3 into the second answer choice?

jimthompson5910 (jim_thompson5910):

what result do you get for y?

OpenStudy (anonymous):

OK.. look at this equation y=4x-10 now if point (3,2) is on this line it must satisfy the equation of line .. satisfying means RHS of equation must be equal to LHS. here on LHS we have y and on RHS we have 4x-10 . putting value of x=3 and y=2 it satisfies the equation so point (3,2) lie on line y=4x-10

OpenStudy (anonymous):

we get y=7 for that X=3

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."